Evaluate 2 a + 4 for each value of a
1.a=0
2.a=2
3.a=5
4.a=8
1
Expert's answer
2020-02-05T09:49:10-0500

We have to find the value of expression "(2a+4)" for some particular values of a :


1). when a=0

"2a+4 = 2*0+4 =0+ 4=4"


2). when a=2

"2a+4 = 2*2+4 = 4+4=8"


3). when a=5

"2a+4 = 2*5+4 =10+ 4=14"


4). when a=8

"2a+4 = 2*8+4 =16+ 4=20"